what is 3-(2c+2)=2(c+2)+3c ?
and you also have to do the three steps
1. do the work
2.check the answer
3.answer

Answer by lolaball(3)

3-(2c+2)=2(c+2)+3c
3-2c-2=2c+4+3c
1-2c=5c+4
+2c=+2c
1=7c+4
-4=-4
-3=7c
-.428??